OpenStudy (anonymous):

**WILL MEDAL** What is the average acceleration of a jet ski that goes from rest to 10 m/s west in 4 s? Take the direction west to be positive, and east to be negative. A. 2.5 m/s2 B. -2.5 m/s2 C. 0.4 m/s2 D. -0.4 m/s2

OpenStudy (anonymous):

Acceleration is rate of change of velocity or (v-u)/t = (10-0)/4= 2.5m /s^2

OpenStudy (anonymous):

So A?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Oh ok thank you thank you thank you!!!

OpenStudy (anonymous):

My pleasure.